### v2

@@ Abstract.
An uncountable $\aleph_{1}$-free group can not admit a Polish group topology but an uncountable $\aleph_{1}$-free abelian group can, as witnessed e.g. by the Baer-Specker group $\mathbb{Z}^{\omega}$, in fact, more strongly, $\mathbb{Z}^{\omega}$ is separable. In this paper we investigate $\aleph_{1}$-free abelian non-Archimedean Polish groups. We prove two main results. The first is that there are continuum many separable (and so torsionless, and so $\aleph_{1}$-free) abelian non-Archimedean Polish groups which are not topologically isomorphic to product groups and are pairwise not continuous homomorphic images of each other. The second is that the following four properties are complete co-analytic subsets of the space of closed abelian subgroups of $S_{\infty}$: separability, torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
@@ 1. Introduction
Recall that a topological group $G$ is said to be Polish if its group topology is Polish, i.e., separable and completely metrizable, and it is said to be non-Archimedean Polish if, in addition, it admits a countable neighborhood basis at the identity consisting of open subgroups. In a meeting in Durham in 1997, D. Evans asked if a non-Archimedean Polish group can be an uncountable free group. Around the same time H. Becker and A. Kechris asked the same question without the assumption of non-Archimedeanity (cf. [1]). In [14] the second author settled the first question, and in the subsequent paper [15] he settled also the second question. It was later discovered that some of the impossibility results from [15] followed already from an old important result of Dudley [5]. All these results were later vastly generalized in [12] by both authors of the present paper in the context of graph product of groups. In another direction, Khelif proved in [10] that no uncountable $\aleph_{1}$-free groups can be Polish, where we recall that a group $G$ is said to be $\aleph_{1}$-free (resp. $\aleph_{1}$-free abelian) if every countable subgroup of $G$ is free (resp. free abelian). Furthermore, a Polish group cannot be free abelian either, by [15]. In contrast, as well-known, the Baer-Specker group $\mathbb{Z}^{\omega}$ when topologized with the product topology is non-Archimedean Polish and $\aleph_{1}$-free (but not free). This interesting state of affairs motivated us to investigate the structure of the class of $\aleph_{1}$-free abelian non-Archimedean Polish groups, wondering in particular: how complicated are these topological groups?
@@
    We $A$ is $\mathbb{Z}$-homogenenous if every element has type $\mathbf{0}$ (cf. 3.7).
@@
[…]
Now, it so happens that $\mathbb{Z}^{\omega}$ is not only $\aleph_{1}$-free but actually separable. Thus, the motivating question of the paper mentioned above (on the existence of $\aleph_{1}$-free abelian non-Archimedean Polish groups) can be strengthen to torsionless (resp. separable) abelian groups. We will deal this all these properties in our paper.
The first result that we present is on the number of groups and, with respect to the four properties mentioned above, it is of the strongest possible form.
@@
There are continuum many separable (hence torsionless, hence $\aleph_{1}$-free) abelian non-Archimedean Polish groups which are not topologically isomorphic to product groups and are pairwise not continuous homomorphic images of each other. Furthermore, all these groups can be taken to be inverse limits of torsion-free completely decomposable groups (i.e., direct sums of $\mathrm{TFAB}$ of rank $1$).
@@
At the heart this program there is the following two-folded task:
@@ 2. $\bullet$
    if a class $\mathcal{C}$ is Borel, study the Borel complexity of topol. isomorphism on $\mathcal{C}$.
@@ 2. The descriptive set theory setting
From here on, as it is customary in set theory, we denote the set of natural numbers by $\omega$. The aim of this section is to fix the coding, i.e., to specify which spaces we refer to in our results, together with introducing the necessary background.
@@
The $G$ be a topological group, then TFAE:
@@
Fact 2.5.
Let $\mathcal{T}(S_{\infty})$ denote the set of trees on $\omega$ with no leaves topologized defining the distance between two trees $T_{1}\neq T_{2}$ as $\frac{1}{2^{n}}$, where $n$ is least $n$ such that $T_{1}\cap\omega^{n}\neq T_{1}\cap\omega^{n}$. For $C\in\mathcal{F}(S_{\infty})$, define:
@@
Then the map $\mathbf{B}$ is a Borel isomorphism from $\mathcal{F}(S_{\infty})$ onto $\mathcal{T}(S_{\infty})$.
In virtue of Fact 2.5, we can work with $\mathcal{T}(S_{\infty})$ instead of $\mathcal{F}(S_{\infty})$.
Fact 2.6.
Proviso 2.7.
From now on we only look at inverse systems with binding onto maps.
Notation 2.8.
We can look at an inverse system $A=(A_{n},f_{(n,m)}:m\leqslant n<\omega)$ as a first-order structure in a language $L_{\mathrm{inv}}=\{P_{n},\cdot_{n},f_{(n,m)}:m\leqslant n<\omega\}$, where:
@@
Notation 2.9.
Of course with respect to the language considered in 2.8 saying that a structure of the form $A=(A_{n},f_{(n,m)}:m\leqslant n<\omega)$ is an inverse system is axiomatizable in first-order logic. Thus, the usual setting of invariant descriptive set theory applies and we can consider the Borel space of countable $\mathrm{Pro}$-$\mathrm{Groups}$ with domain (a subset of) $\omega$. For concreteness we also assume that our inverse systems $A=(A_{n},f_{(n,m)}:m\leqslant n<\omega)$ are such that e.g. $\mathrm{dom}(A_{n})\subseteq\{p_{n}^{m}:0<m<\omega\}$, where $p_{n}$ is the $n$-th prime number, so that the different sorts $A_{n}$’s are disjoint, and thus in particular $n\neq m$ implies that $e_{A_{n}}\neq e_{A_{m}}$ (where $e_{A_{n}}$ is the neutral element of $A_{n}$). Finally, naturally when the groups $A_{n}$ are abelian, then we move to additive notation and so we write $+_{n}$ instead of $\cdot_{n}$ and $0_{A_{n}}$ instead of $e_{A_{n}}$.
@@
$\pi_{\bar{a}}(k)=\begin{cases}k+_{A_{n}}a_{n},\;\;\,\text{ if there is }n<\omega\text{ s.t. }k\in A_{n},\\ k,\;\;\;\;\;\;\;\;\;\;\;\;\;\;\,\,\text{ if }k\notin\bigcup_{n<\omega}A_{n}.\end{cases}$
Observation 2.11.
@@
We omit the details of the next proposition as its proof is standard (use e.g. 2.5).
@@ 3.1. The engine
The main source of construction of Polish non-Archimedean $\mathrm{TFAB}$ present in this paper is described in the following definition, which we refer to as the engine.
@@
Let $T\subseteq\omega^{<\omega}$ be a tree with no leaves and $L:T\rightarrow\mathcal{P}(\mathbb{P})$ be s.t.:
@@
    for $m\leqslant n<\omega$, $f_{(n,m)}$ is the unique homomorphism from $G_{n}$ into $G_{m}$ s.t.:
    $\eta\in T\cap\omega^{n}\;\Rightarrow\;x_{\eta}\mapsto x_{\eta\restriction m}.$
Notice that the inverse system $\mathrm{inv}(T,L)=(G_{n},f_{(n,m)}:m\leqslant n<\omega)$ is well-defined because of conditions (a) and (b) above. Being it well-defined, we do the following:
@@
Notice that the inverse systems $\mathrm{inv}(T,L)=(G_{n},f_{(n,m)}:m\leqslant n<\omega)$ defined in 3.2 have bounding maps $f_{(n,m)}$ which are onto; this is because we ask that the tree $T$ has no leaves and because we ask that condition (b) from there holds.
@@
Notice also that for $\mathrm{inv}(T,L)=(G_{n},f_{(n,m)}:m\leqslant n<\omega)$ as in 3.2 we have that the groups $G_{n}$ are completely decomposable, i.e., they are direct sums of groups of rank $1$, in fact it follows immediately from the definitions that:
@@
Proof.
This is immediate. ∎
Notation 3.5.
Given a tree $T\subseteq\omega^{<\omega}$ we denote by $[T]$ the set of $\eta\in\omega^{\omega}$ such that $\eta\restriction n\in T$ for all $n<\omega$, i.e., the set of infinite branches of $T$. Recall that in 3.2 we only consider trees with no leaves and so, for $T$’s as in 3.2, $[T]$ is always non-empty.
For background on the notions that we now introduce see e.g. [7, Chapter 13].
@@
Let $A\in\mathrm{TFAB}$ and let $(p_{i}:i<\omega)$ be the list of the prime numbers in increasing order. For $a\in A$, we define the characteristic of $a$, denoted as $\chi(a)$, as follows: $\chi(a)=(h_{p_{i}}(a):i<\omega),$ where $h_{p_{i}}(a)$ is the supremum of the $k<\omega$ such that $p_{i}^{k}\,|\,a$, where the supremum is taken in $\omega\cup\{\infty\}$ (so the value $\infty$ is allowed).
@@
    We denote by $\mathbf{0}$ the type of the characteristic $(0,0,...)$.
@@
Fact 3.8 ([7, (C) on pg. 411]).
Let $G$ and $H$ be torsion-free abelian and suppose that there is an homomorphism $f:G\rightarrow H$, then, for every $g\in G$, $\mathbf{t}(g)\leqslant\mathbf{t}(f(g))$.
Fact 3.9.
@@ 3.2. Nice pairs
    $S\subseteq T$ and $\eta\in S$ implies that $0\notin\mathrm{ran}(\eta)$;
@@
    $L$ is the function with domain $T$ defined as follows:
    $L(\eta)=\{p_{\nu}:(\nu\triangleleft\eta\text{ or }\eta\trianglelefteqslant\nu)\text{ and }\nu\in S^{+}\}.$
Observation 3.11.
@@
Convention 3.12.
For the rest of this subsection, we fix $(T,L)$, $S$, $S^{+}$ and $\bar{p}$ as in 3.10 and we let $G=G(T,L)$. For $\eta\in[T]$, we let $L(\eta)=\bigcap\{L(\eta\restriction n):n<\omega\}$. Also, for $n<\omega$, we let $f_{(\omega,n)}$ be the obvious projection of $G$ onto $G_{n}$. For $\nu\in[T]$, we let $x_{\nu}=(x_{\nu\restriction n}:n<\omega)\in G$.
@@
(Recalling 3.12.) If $S$ has an $\omega$-branch (i.e., there is $\eta\in[T]$ such that, for all $n<\omega$, $\eta\restriction n\in S$), then $G$ is not $\mathbb{Z}$-homogeneous, and so $G$ is not $\aleph_{1}$-free.
@@
For $\eta\in T\cup[T]$, let $K_{\eta}$ be the additive subgroup of $\mathbb{Q}$ generated by $K^{-}_{\eta}=\{\frac{1}{p}:p\in L(\eta)\}\subseteq\mathbb{Q}$, explicitly, $K_{\eta}$ is the subgroup of $\mathbb{Q}$ consisting of:
$\{\frac{b}{\prod_{\ell<k}p_{\ell}}:b\in\mathbb{Z},\;k<\omega,\;p_{\ell}\in L(\eta),\;(p_{\ell}:\ell<k)\text{ with no repetitions}\}.$
@@
If $S$ has no $\omega$-branch, then $G$ is separable (and so $G$ torsionless, and so $G$ is $\aleph_{1}$-free).
@@
We first show that $G$ is torsionless. For any $y\in G\setminus\{0\}$ we want to find $f\in\mathrm{Hom}(G,\mathbb{Z})$ such that $f(y)\neq 0$. To this extent, fix $y\in G\setminus\{0\}$, so $y=(y_{n}:n<\omega)$.
  1. $(\star_{1})$
    For $n<\omega$, let $y_{n}=\sum\{a_{\nu}x_{\nu}:\nu\in Y_{n}\}$, where:
    1. $(\cdot_{1})$
      $Y_{n}\subseteq\omega^{n}\cap T$ is finite;
    2. $(\cdot_{2})$
      $a_{\nu}\in K_{\nu}\setminus\{0\}$ (recall 3.14).
  1. $(\star_{2})$
    We can find $n_{*}<\omega$ such that, for every $n_{*}\leqslant n<\omega$, $y_{n}=f_{(\omega,n)}(y)\neq 0$.
  1. $(\star_{3})$
    By induction on $n\geqslant n_{*}$ we choose $\nu_{n}$ and $b_{\nu}$ such that:
@@ 1. $(\star_{3})$
      $\nu_{n}\in Y_{n}$;
    2. (b)
      $n=m+1>n_{*}$ implies $\nu_{m}\triangleleft\nu_{n}$;
Why can we carry the induction? For $n=n_{*}$ let $\nu_{n}\in Y_{n}$. This is possible by $(\star_{2})$. For $n=m+1$, note that $f_{({n}+1,k_{n})}\circ f_{(\omega,{n}+1)}=f_{(\omega,{n})}$ and so $a_{\nu_{m}}=\sum\{a_{\rho}:\nu_{m}\triangleleft\rho\in Y_{n+1}\}$. As $\nu_{m}\in Y_{m}$ we have $a_{\nu_{n}}\neq 0$, hence there is $\rho\in Y_{n+1}$ such that $\nu_{m}\triangleleft\rho$, so letting $\nu_{n}=\rho$ we can satisfy clauses (a) and (b).
  1. $(\star_{4})$
    Let $\nu=\bigcup\{\nu_{n}:n\geqslant n_{*}\}\in[T]$, for $(\nu_{n}:n\geqslant n_{*})$ as in $(\star_{3})$.
Now there are two cases.
Case 1. For all $n<\omega$, $\nu\restriction n\in S$.
This is impossible, as we are assuming that $S$ has no $\omega$-branch.
Case 2. For some $n<\omega$, $\nu\restriction n\notin S$.
If this is the case, then w.l.o.g. $n>n_{*}$ and easily we have that $L(\nu\restriction(n+1))$ is finite. Let $\eta=\nu_{n+1}$.
  1. $(\star_{5})$
    Let $g\in\mathrm{Hom}(G,K_{\eta}x_{\eta})$ be defined for $z\in G$ by $g(z)=\pi_{(n,\eta)}(f_{(\omega,n)})(z)$, where $\pi_{(n,\eta)}$ if the projection from $G_{\mathrm{lg}(\eta)}$ onto $\langle x_{\nu}\rangle^{*}_{G_{\mathrm{lg}}(\eta)}$ (recall that $G_{\mathrm{lg}(\eta)}$ is completely decomposable, cf. 3.4).
Clearly we have that:
  1. $(\star_{6})$
    1. (a)
      $g(y)=a_{\eta}x_{\nu}\neq 0$;
    2. (b)
      $g\in\mathrm{Hom}(G,K_{\eta}x_{\eta})$.
[Why? See $(\star_{1})(\cdot_{2})$.]
But then, as $L(\eta)$ is finite, clearly $K_{\eta}x_{\eta}\cong\mathbb{Z}x_{\eta}$, and so we are done proving that $G$ is torsionless.
We now show that $G$ is actually separable. To this extent, let $X\subseteq G$ be finite.
  1. $(*_{1})$
    It suffices to prove that there are $n<\omega$ and finite $Y\subseteq\omega^{n}\cap T$ such that letting $\pi_{(n,Y)}$ be the natural projection of $G_{n}$ onto $G_{(n,Y)}:=\langle Y\rangle^{*}_{G_{n}}$ we have:
    1. (a)
      for every $\nu\in Y$ we have that $L(\nu)$ is finite;
@@ 1. $(*_{1})$
We show that $(*_{1})$ holds. To this extent, first of all, notice that $G_{(n,Y)}$ is a free direct summand of $G$, because $G_{n}=\bigoplus\{K_{\nu}x_{\nu}:\nu\in T\cap\omega\}$ and, for every $\nu\in Y$, $L(\nu)$ is finite and so $K_{\nu}\cong\mathbb{Z}$. Secondly, observe that then also $\langle X\rangle^{*}_{G}$ is free, since by (c), $g_{0}$ is an isomorphism. Observe now that, by clause (b), we have that $g_{0}^{-1}\circ g$ is an homomorphism from $G$ into $\langle X\rangle^{*}_{G}$. Also, by clause (c), clearly $g_{0}^{-1}$ is an isomorphism from $G_{(n,Y)}$ onto $\langle X\rangle^{*}_{G}$. Notice now that for $a\in\langle X\rangle^{*}_{G}$ we have that:
@@
Hence, we have that $g_{0}^{-1}\circ g$ is a retraction of $G$ onto the free subgroup $\langle X\rangle^{*}_{G}$ and so $X$ is contained in a free direct summand of $G$, as desired.

### v3

@@ Abstract.
An uncountable $\aleph_{1}$-free group cannot admit a Polish group topology but an uncountable $\aleph_{1}$-free abelian group can, as witnessed, for example, by the Baer-Specker group $\mathbb{Z}^{\omega}$; more strongly, $\mathbb{Z}^{\omega}$ is separable. In this paper we investigate $\aleph_{1}$-free abelian non-Archimedean Polish groups. We prove two main results. The first is that there are continuum many separable (and so torsionless, and so $\aleph_{1}$-free) abelian non-Archimedean Polish groups which are pairwise not topologically isomorphic. The second is that the following four properties are complete co-analytic subsets of the space of closed abelian subgroups of $S_{\infty}$: separability, torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
@@ 1. Introduction
Recall that a topological group $G$ is said to be Polish if its group topology is Polish, i.e., separable and completely metrizable, and it is said to be non-Archimedean Polish if, in addition, it admits a countable neighborhood basis at the identity consisting of open subgroups. In a meeting in Durham in 1997, D. Evans asked if a non-Archimedean Polish group can be an uncountable free group. Around the same time H. Becker and A. Kechris asked the same question without the assumption of non-Archimedeanity (cf. [1]). In [14] the second author settled the first question, and in the subsequent paper [15] he settled also the second question. It was later discovered that some of the impossibility results from [15] followed already from an old important result of Dudley [5]. All these results were later vastly generalized in [12] by both authors of the present paper in the context of graph products of groups. In another direction, Khelif proved in [10] that no uncountable $\aleph_{1}$-free groups can be Polish, where we recall that a group $G$ is said to be $\aleph_{1}$-free (or $\aleph_{1}$-free abelian) if every countable subgroup of $G$ is free (or free abelian, respectively). Furthermore, an uncountable Polish group cannot be free abelian either, by [15]. In contrast, as well-known, the Baer-Specker group $\mathbb{Z}^{\omega}$ when topologized with the product topology is non-Archimedean Polish and $\aleph_{1}$-free abelian (but not free abelian). This interesting state of affairs motivated us to investigate the structure of the class of $\aleph_{1}$-free abelian non-Archimedean Polish groups, wondering in particular: how complicated is the class of these topological groups?
@@
    We say that $A$ is $\mathbb{Z}$-homogenenous if every element has type $\mathbf{0}$ (cf. 3.6).
@@
$\text{separable }\Rightarrow\text{ torsionless }\Rightarrow\text{ $\aleph_{1}$-free }\Rightarrow\text{$\mathbb{Z}$-homogenenous}.$
Now, it so happens that $\mathbb{Z}^{\omega}$ is not only $\aleph_{1}$-free but actually separable. Thus, the motivating question of the paper mentioned above (on the existence of $\aleph_{1}$-free abelian non-Archimedean Polish groups) can be strengthened to torsionless, or even separable, abelian groups. We will deal with all these properties in our paper.
The first result concerns the number of non-isomorphic groups realizing the four properties above, and it is of the strongest possible form. We recall that a completely decomposable abelian group is a group that is a direct sum of groups of rank $1$.
@@
There are continuum many separable (hence torsionless, hence $\aleph_{1}$-free) abelian non-Archimedean Polish groups which are pairwise not topologically isomorphic. Furthermore, all these groups can be taken to be inverse limits of countable torsion-free completely decomposable groups.
@@
At the heart this program there is the following two-fold task:
@@ 2. $\bullet$
    if a class $\mathcal{C}$ is Borel, study the Borel complexity of topological isomorphism on $\mathcal{C}$.
@@
The structure of the paper is as follows. In Section 2 we collect the necessary preliminaries. In Section 3 we introduce the main group-theoretic construction at the heart of the paper. In Section 4 we prove Theorems 1.2 and 1.3. Finally, in Section 5 we establish some complementary results.
@@ 2. The descriptive set theory setting
From here on, as is customary in set theory, we denote the set of natural numbers by $\omega$. The aim of this section is to fix the coding, i.e., to specify which spaces we refer to in our results, together with introducing the necessary background.
@@
Let $G$ be a topological group, then the following are equivalent:
@@
Notation 2.6.
Given a tree $T\subseteq\omega^{<\omega}$ we denote by $[T]$ the set of $\eta\in\omega^{\omega}$ such that $\eta\restriction n\in T$ for all $n<\omega$, i.e., the set of infinite branches of $T$.
Fact 2.7.
Let $\mathcal{T}(S_{\infty})$ denote the set of trees on $\omega$ with no leaves topologized defining the distance between two trees $T_{1}\neq T_{2}$ as $\frac{1}{2^{n}}$, where $n$ is least $n$ such that $T_{1}\cap\omega^{n}\neq T_{2}\cap\omega^{n}$. For $C\in\mathrm{Sgp}(S_{\infty})$, define $T_{C}\in\mathcal{T}(S_{\infty})$ as follows:
@@
Then $\mathbf{B}:\mathrm{Sgp}(S_{\infty})\rightarrow\mathcal{T}(S_{\infty})$ is Borel, it is injective and it has closed range; furthermore, when restricted to its range, the map $\mathbf{B}$ is a Borel isomorphism.
By Fact 2.7, we can work with $\mathbf{B}(\mathrm{Sgp}(S_{\infty}))\subseteq\mathcal{T}(S_{\infty})$ instead of $\mathrm{Sgp}(S_{\infty})$.
Proviso 2.8.
From now on we only look at inverse systems with onto bonding maps.
Notation 2.9.
We can look at an inverse system $A=(A_{n},f_{(n,m)}:m\leqslant n<\omega)$ of groups as a first-order structure in a language $L_{\mathrm{inv}}=\{P_{n},\cdot_{n},f_{(n,m)}:m\leqslant n<\omega\}$, where:
@@
Notation 2.10.
Of course with respect to the language considered in 2.9 saying that a structure of the form $A=(A_{n},f_{(n,m)}:m\leqslant n<\omega)$ is an inverse system is axiomatizable in first-order logic. Thus, we can consider the Borel space of countable $\mathrm{Pro}$-$\mathrm{Groups}$ with domain (a subset of) $\omega$. We also assume that our inverse systems $A=(A_{n},f_{(n,m)}:m\leqslant n<\omega)$ are such that for $n\neq m$, $\mathrm{dom}(A_{n})\cap\mathrm{dom}(A_{m})=\emptyset$. Finally, naturally when the groups $A_{n}$ are abelian, then we move to additive notation and so we write $_{n}$ instead of $\cdot_{n}$ and $0_{A_{n}}$ instead of $e_{A_{n}}$.
@@
Observation 2.12.
[…]
We omit the details of the next proposition as its proof is standard (use e.g. 2.7).
@@ 3.1. The engine
The following construction will be used to prove Theorems 1.2 and 1.3.
@@
Let $T\subseteq\omega^{<\omega}$ be a tree with no leaves and $L:T\rightarrow\mathcal{P}(\mathbb{P})$ be such that:
@@
    for $m\leqslant n<\omega$, $f_{(n,m)}$ is the unique homomorphism from $G_{n}$ into $G_{m}$ s.t.: η∈T ∩ω^n   ⇒  x_η↦x_η↾m.
Notice that the inverse system $\mathrm{inv}(T,L)=(G_{n},f_{(n,m)}:m\leqslant n<\omega)$ is well-defined because of conditions (a) and (b) above. It being well-defined, we do the following:
@@
Notice that the inverse systems $\mathrm{inv}(T,L)=(G_{n},f_{(n,m)}:m\leqslant n<\omega)$ defined in 3.2 have bonding maps $f_{(n,m)}$ which are onto; this is because we ask that the tree $T$ has no leaves and because we ask that condition (b) from there holds. Recall that in 3.2 we only consider trees with no leaves and so, for $T$’s as in 3.2, the set $[T]$ is always non-empty (cf. 2.6 for a definition of $[T]$).
@@
Notice also that for $\mathrm{inv}(T,L)=(G_{n},f_{(n,m)}:m\leqslant n<\omega)$ as in 3.2 we have that the groups $G_{n}$ are completely decomposable, i.e., they are direct sums of groups of rank 11, in fact it follows immediately from the definitions that:
@@
For background on the notions that we now introduce see e.g. [7, Chapter 13]. We will often abbreviate “torsion-free abelian group” with TFAB.
@@
Let $A\in\mathrm{TFAB}$ and let $(p_{i}:i<\omega)$ be the list of the prime numbers in increasing order. For $a\in A$, we define the characteristic of $a$ in $A$, denoted as $\chi_{A}(a)=\chi(a)$, as follows: $\chi(a)=(h_{p_{i}}(a):i<\omega),$ where $h_{p_{i}}(a)$ is the supremum of the $k<\omega$ such that $p_{i}^{k}\,|\,a$, where the supremum is taken in $\omega\cup\{\infty\}$ (so the value $\infty$ is allowed).
@@
    We denote by $\mathbf{0}$ the type of the characteristic (0,0,)($0,$...).
@@
Fact 3.7 ([7, (C) on pg. 411]).
Let $G$ and $H$ be torsion-free abelian and suppose that there is a homomorphism $f:G\rightarrow H$, then, for every $g\in G$, $\mathbf{t}(g)\leqslant\mathbf{t}(f(g))$.
Fact 3.8.
@@ 3.2. Nice pairs
We now present a technical definition necessary for our approach to work, the fact that this or similar technical adjustments are necessary is justified by Section 5.2.
@@
    $\eta\in S$ implies that $0\notin\mathrm{ran}(\eta)$;
@@
    $L$ is the function with domain $T$ defined as follows: L(η) = { p_ν: (νη or ην) and ν∈S^+}.
Observation 3.10.
@@
Convention 3.11.
For the rest of this subsection, we fix $(T,L)$, $S$, $S^{+}$ and $\bar{p}$ as in 3.9 and we let $G=G(T,L)$. For $\eta\in[T]$, we let $L(\eta)=\bigcap\{L(\eta\restriction n):n<\omega\}$. Also, for $n<\omega$, we let $f_{(\omega,n)}$ be the obvious projection of $G$ onto $G_{n}$. For $\nu\in[T]$, we let $x_{\nu}=(x_{\nu\restriction n}:n<\omega)\in G$. Recall also that $f_{(n,m)}$ denote the bonding map from $G_{n}$ onto $G_{m}$.
@@
If $S$ has an $\omega$-branch (i.e., there is $\eta\in[T]$ such that, for all $n<\omega$, $\eta\restriction n\in S$), then $G$ is not $\mathbb{Z}$-homogeneous, and so $G$ is not $\aleph_{1}$-free.
@@
For $\eta\in T\cup[T]$, let $K_{\eta}$ be the additive subgroup of $\mathbb{Q}$ generated by $\{\frac{1}{p}:p\in L(\eta)\}\subseteq\mathbb{Q}$; explicitly, $K_{\eta}$ is the subgroup of $\mathbb{Q}$ consisting of:
$\{\frac{b}{\prod_{\ell<k}p_{\ell}}:b\in\mathbb{Z},\;k<\omega,\;p_{\ell}\in L(\eta)\}.$
@@
If $S$ has no $\omega$-branch, then $G=G(T,L)$ is separable (and so $G$ is torsionless, and so $G$ is $\aleph_{1}$-free).
@@
We show that $G$ is separable. To this end, let $X\subseteq G$ be finite.
  1. $(*_{1})$
    We claim that it suffices to prove that there are $n<\omega$ and finite $Y\subseteq\omega^{n}\cap T$ such that letting $\pi_{(n,Y)}$ be the natural projection of $G_{n}$ onto $G_{(n,Y)}:=\langle Y\rangle^{*}_{G_{n}}$ we have:
@@ 1. $(*_{1})$
      for every $\nu\in Y$, $L(\nu)$ is finite;
@@
We show that $(*_{1})$ holds, i.e., that the conditions listed there suffice to conclude that $G$ is separable. For this purpose, first of all, notice that $G_{(n,Y)}$ is a free direct summand of $G$, because $G_{n}=\bigoplus\{K_{\nu}x_{\nu}:\nu\in T\cap\omega^{n}\}$ and, for every $\nu\in Y$, $L(\nu)$ is finite and so $K_{\nu}\cong\mathbb{Z}$. Secondly, observe that then also $\langle X\rangle^{*}_{G}$ is free, since by (c), $g_{0}$ is an isomorphism. Observe now that, by clause (b), we have that $g_{0}^{-1}\circ g$ is an homomorphism from $G$ into $\langle X\rangle^{*}_{G}$. Also, by clause (c), clearly $g_{0}^{-1}$ is an isomorphism from $G_{(n,Y)}$ onto $\langle X\rangle^{*}_{G}$. Notice now that for $a\in\langle X\rangle^{*}_{G}$ we have that:
@@
Hence, we have that $g_{0}^{-1}\circ g$ is a retraction of $G$ onto the free subgroup $\langle X\rangle^{*}_{G}$ and so $X$ is contained in a free direct summand of $G$, as desired. Hence, $(*_{1})$ holds indeed.
So now that we have showed that $(*_{1})$ suffices, for the rest of the proof we will focus on showing that the conditions from $(*_{1})$ can be fulfilled (cf. the final $(*_{19})$).